# 2025-04-30

### v1

@@ 5 Numerical Results
We demonstrate the applicability of the SDE-AL framework over two parameter-dependent dynamical systems. The first system corresponds to a problem setup featuring fluid flow behavior governed by the shallow water equations, which is challenging due to the numerous shock interactions across the spatial domain as the system evolves in time. The nature of these interactions is affected by the fluid viscosity. Next, we consider a problem showcasing a fluid flow with a vortex shedding phenomena governed by the incompressible Navier-Stokes equations. Here, the shedding behavior is dependent on the Reynolds number under consideration.
For brevity, we present the results from the SDE-ActLearn-POD-KSNN ROM for the shallow water equations, whereas, the results from the SDE-Actearn-POD-NN ROM are presented for the Navier-Stokes equations. Moreover, variant-A of the SDE-AL framework, i.e., Algorithm 3.2, is tested for the shallow water equations example, whereas, variant-B of the SDE-AL framework, i.e., Algorithm 3.1, is tested for the Navier-Stokes equations example. Furthermore, for the shallow water equations example, we also provide a comparison between the SDE-AL results obtained while using $\mathcal{D}_{1}$ and those while using $\hat{\mathcal{D}}_{2}$. This comparison highlights the benefit of using the newly proposed distance metric for performing SDE-AL.
@@ 5.1 Shallow Water Equations
$\partial_{t}h+\partial_{x}(hu)$ $=0,$ (5.1)
$\partial_{t}(hu)+\partial_{x}\Big{(}hu^{2}+\frac{1}{2}gh^{2}\Big{)}$ $=-\frac{\nu}{\lambda}u,$ (5.2)
@@
$h(0,x)$ $=1+\exp(3\cos(\pi(x+0.5))-4).$ (5.3)
$u(0,x)$ $=0.25.$ (5.4)
The problem setup and FOM solver are kept similar to [32], allowing us to easily compare the results of the two active learning approaches. More specifically, we fix $\lambda=0.1$, and consider $\nu$ as the free parameter, ranging from $1$ to $10^{-5}$. Furthermore, we consider $100$ viscosity values in total, accounting for both the parameter sets $P$ and $P^{*}$. The parameter samples considered during active learning are $100$ evenly distributed logarithmic $\nu$ values in $[\log_{10}(10^{-5}),\log_{10}(1)]$. The spatial domain $\Omega\in[-1,1]$ has $601$ grid nodes and is with periodic boundary conditions. FOM solutions corresponding to each training parameter sample are collected at $200$ uniform time steps in $t\in[0,2]$, creating parameter-specific snapshot matrices.
@@
The distances between all neighboring parameter-specific subspaces are evaluated and the parameter pair with the farthest subspaces is identified, followed by the selection of a new sample from the candidate set $P^{*}$ as outlined in Algorithm 3.2. This process is continued iteratively by updating the neighboring parameter pairs, evaluating the distance between new pairs, and selecting new samples from $P^{*}$ in parameter regions surrounded by the subspace pair that are farthest.
We compare two active learning procedures following Algorithm 3.2, by using $\mathcal{D}_{1}$ and $\hat{\mathcal{D}}_{2}$. While using $\mathcal{D}_{1}$, the variation of maximal distance value $\mathcal{D}_{max}^{(iter)}$ with iterations is quite erratic and takes $31$ parameter selections to attain a distance tolerance $tol=0.2$, and an error tolerance $tol=10^{-2}$. Whereas, with distance criterion $\hat{\mathcal{D}}_{2}$, the maximal distance is seen to monotonically decrease, also at a faster rate. Moreover, we notice that the SDE-AL procedure employing $\hat{\mathcal{D}}_{2}$ converges with just $11$ new parameter selections.
@@
The SDE-ActLearn-POD-KSNN ROMs are queried at the following out-of-training viscosity values: $\{5\times 10^{-1},5\times 10^{-2},5\times 10^{-3},5\times 10^{-4},5\times 10^{-5}\}$. By querying the ROM constructed using the distance criterion $\hat{\mathcal{D}}_{2}$, Figures 5.3 and 5.4 show the fluid height and velocity predictions at three representative test samples, along with their respective relative solution errors, in the entire space-time domain. The reported relative error values are the absolute of the difference between the FOM and ROM solutions, normalized by the $2$-norm of the FOM solution. It is clear from Figures 5.3 and 5.4 that the ROM is able to generate very accurate solution profiles, successfully capturing the multiple shock interactions over time. A similar solution accuracy is observed for the SDE-ActLearn-POD-KSNN ROM constructed using the criterion $\mathcal{D}_{1}$. As a result, to avoid similar figures, we do not show the solution and error contour plots for it.
@@ 5.2 Incompressible Navier-Stokes Equations
The incompressible Navier-Stokes equations are used to model the behavior of viscous fluids. The following form of $2D$ Navier-Stokes equations is considered:
$\rho\left(\partial_{t}\mathbf{u}+\mathbf{u}\cdot\nabla\mathbf{u}\right)$ $=\mu\nabla\cdot\left(\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}\right)-\nabla p,$ (5.6)
$\nabla\cdot\mathbf{u}$ $=0,$ (5.7)
where $\mathbf{u}(t,\mathbf{z})\in\mathbb{R}^{2}$ is the velocity vector and $p(t,\mathbf{z})\in\mathbb{R}$ is the fluid pressure, with time $t\geq 0$ and space $\mathbf{z}:=(x\times y)\in\Omega\in\mathbb{R}^{2}$. The density and dynamic viscosity of the fluid are denoted by $\rho$ and $\mu$, respectively.
@@
$\mathbf{u}$ $=\left[\frac{4Uy(0.41-y)}{0.41^{2}},0\right]^{\top},$ $\mathbf{z}\in\Gamma_{in},$ (5.8)
$\mu\left(\nabla\mathbf{u}+(\nabla\mathbf{u})^{\top}\right)\cdot\mathbf{n}$ $=p\mathbf{n},$ $\mathbf{z}\in\Gamma_{out},$ (5.9)
$\mathbf{u}$ $=0,$ $\mathbf{z}\in\Gamma_{cyl}\cup\Gamma_{top}\cup\Gamma_{bottom},$ (5.10)
@@
The domain $\Omega_{fluid}$ is discretized with tetrahedral mesh elements amounting to a total of $37,514$ grid nodes. The Navier-Stokes equations are numerically solved using the FEniCS package [42]. For further details about the implementation, please refer [32]. We fix the fluid density to a value of $\rho=1\ kg/m^{3}$ and can freely vary the dynamic viscosity $\mu$, resulting in a variable Reynolds number between $[80,200]$. In this work, we focus on constructing an actively-learned ROM for the fluid velocity magnitude over the complete domain $\Omega_{fluid}$.
To begin the active learning procedure, $81$ evenly distributed $Re$ values are considered from $[80,200]$. Out of these, five $Re$ values, i.e., $\{80,110,140,170,200\}$, are initially taken in the parameter set $P$, whereas, the remaining $76$ $Re$ values are placed in the candidate set $P^{*}$. Furthermore, parameter-specific snapshot matrices for the fluid magnitude are constructed by querying the FOM solver corresponding to each of the $Re$ value from set $P$. Similar to the shallow water equations example, parameter-specific subspaces are then formed, retaining $99.9999\%$ of the total energy, which corresponds to $\eta=10^{-6}$.
@@
With this problem, we intend to demonstrate the applicability of Algorithm 3.1 and terminate the active learning procedure as soon as a specific user-defined computational budget for FOM queries is satisfied. Moreover, the SDE-ActLearn-POD-NN ROM is constructed as outlined in Algorithm 4.7. Once the initial set of parameter-specific subspaces are formed, the distances between all the neighboring subspaces are evaluated and the parameter pair with the farthest subspaces is identified. This is followed by the selection of a new sample from the candidate set $P^{*}$ as outlined in Algorithm 3.1. In an iterative fashion, this procedure is continued, by updating the neighboring parameter pairs, evaluating the distance between new pairs, and selecting new samples from $P^{*}$ in parameter regions surrounded by subspaces that are farthest.
@@
Table 5.2 provides runtime for the FOM solver and the SDE-ActLearn-POD-NN ROM. The FOM solver is executed on a node of the compute cluster Mechthild available at the Max Planck Institute Magdeburg. The node is equipped with two Intel® Xeon Silver 4110 central processing units and $192$ GB of RAM. The deep NN is trained on an NVIDIA Tesla P100 graphics processing unit associated with a similar cluster node. Like the shallow water equations example, we report the time required for creating the parameter-specific subspaces along with all the distance evaluations during the active learning process under SDE-AL in Table 5.2. This amounts to $59.8689$ seconds. The runtime reported under ‘Remaining’ corresponds to the global reduced basis construction in eqs. 4.16 and 4.17, NN data preparation, as well as the NN model construction and training. Out of this, the network training takes $2072.5892$ seconds. Overall, the offline phase cost is dominated by computations towards the FOM queries. The SDE-AL framework enables selection of critical parameter locations in the offline phase, thereby, rendering an effective ROM construction. In comparison with a FOM solver query, the ROM’s online query time of $0.0043$ seconds provides us a computational speedup of greater than $\mathcal{O}(10^{6}).$

### v2

@@ 5 Numerical Results
We demonstrate the applicability of the SDE-AL framework for two parameter-dependent dynamical systems. The first system corresponds to a problem setup featuring fluid flow behavior governed by the shallow water equations, which is challenging due to the numerous shock interactions across the spatial domain as the system evolves in time. The nature of these interactions is affected by the fluid viscosity. Next, we consider a problem showcasing a fluid flow with a vortex shedding phenomena governed by the incompressible Navier-Stokes equations. Here, the shedding behavior depends upon the Reynolds number under consideration.
For brevity, we present the results from the SDE-ActLearn-POD-KSNN ROM for the shallow water equations, whereas, the results from the SDE-ActLearn-POD-NN ROM are presented for the Navier-Stokes equations. Moreover, variant-A of the SDE-AL framework, i.e., Algorithm 3.2, is tested for the shallow water equations example, whereas, variant-B of the SDE-AL framework, i.e., Algorithm 3.1, is tested for the Navier-Stokes equations example. Furthermore, for the shallow water equations example, we also provide a comparison between the SDE-AL results obtained while using $\mathcal{D}_{1}$ and those while using $\hat{\mathcal{D}}_{2}$. This comparison highlights the benefit of using the newly proposed distance metric for performing SDE-AL.
@@ 5.1 Shallow Water Equations
$\partial_{t}h+\partial_{x}(hu)$ $=0,$
$\partial_{t}(hu)+\partial_{x}\Big{(}hu^{2}+\frac{1}{2}gh^{2}\Big{)}$ $=-\frac{\nu}{\lambda}u,$
@@
$h(0,x)$ $=1+\exp(3\cos(\pi(x+0.5))-4),$
$u(0,x)$ $=0.25.$
The problem setup and the FOM solver are kept similar to [32], allowing us to easily compare the results of the two active learning approaches. The system of equations is solved using the discontinuous Galerkin method, where the local Lax–Friedrichs Riemann solver is employed to evaluate the numerical flux, and a second-order strong stability preserving Runge–Kutta scheme is used for time integration. We fix $\lambda=0.1$ and consider $\nu$ as the free parameter, ranging from $1$ to $10^{-5}$. Furthermore, we consider $100$ viscosity values in total, accounting for both the parameter sets $P$ and $P^{*}$. The parameter samples considered during active learning are $100$ evenly distributed logarithmic $\nu$ values in $[\log_{10}(10^{-5}),\log_{10}(1)]$. The spatial domain $\Omega\in[-1,1]$ has $601$ grid nodes and is equipped with periodic boundary conditions. FOM solutions corresponding to each training parameter sample are collected at $200$ uniform time steps in $t\in[0,2]$, creating parameter-specific snapshot matrices.
@@
The distances between all neighboring parameter-specific subspaces are evaluated and the parameter pair with the most distant subspaces is identified, followed by the selection of a new sample from the candidate set $P^{*}$ as outlined in Algorithm 3.2. This process is continued iteratively by updating the neighboring parameter pairs, evaluating the distance between new pairs, and selecting new samples from $P^{*}$ in parameter regions surrounded by the subspace pair that are farthest away.
We compare two active learning procedures following Algorithm 3.2, by using $\mathcal{D}_{1}$ and $\hat{\mathcal{D}}_{2}$. When using $\mathcal{D}_{1}$, the variation of maximal distance value $\mathcal{D}_{max}^{(iter)}$ with iterations is quite erratic and takes $31$ parameter selections to attain a distance tolerance $tol=0.2$, and an error tolerance $tol=10^{-2}$. Whereas, with distance criterion $\hat{\mathcal{D}}_{2}$, the maximal distance is seen to monotonically decrease, also at a faster rate. Moreover, we notice that the SDE-AL procedure employing $\hat{\mathcal{D}}_{2}$ converges with just $11$ new parameter selections.
@@
The SDE-ActLearn-POD-KSNN ROMs are queried at the following out-of-training viscosity values: $\{5\times 10^{-1},5\times 10^{-2},5\times 10^{-3},5\times 10^{-4},5\times 10^{-5}\}$. By querying the ROM constructed using the distance criterion $\hat{\mathcal{D}}_{2}$, Figures 5.3 and 5.4 show the fluid height and velocity predictions at three representative test samples, along with their respective relative solution errors, in the entire space-time domain. The reported relative error values are the absolute difference between the FOM and ROM solutions, normalized by the $2$-norm of the FOM solution. It is clear from Figures 5.3 and 5.4 that the ROM is able to generate very accurate solution profiles, successfully capturing the multiple shock interactions over time. A similar solution accuracy is observed for the SDE-ActLearn-POD-KSNN ROM constructed using the criterion $\mathcal{D}_{1}$. As a result, to avoid similar figures, we do not show the solution and error contour plots for it.
@@ 5.2 Incompressible Navier-Stokes Equations
The incompressible Navier-Stokes equations are used to model the behavior of viscous fluids. The following form of the $2D$ Navier-Stokes equations is considered:
$\rho\left(\partial_{t}\mathbf{u}+\mathbf{u}\cdot\nabla\mathbf{u}\right)$ $=\mu\nabla\cdot\left(\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}\right)-\nabla p,$
$\nabla\cdot\mathbf{u}$ $=0,$
where $\mathbf{u}(t,\mathbf{z})\in\mathbb{R}^{2}$ is the velocity vector and $p(t,\mathbf{z})\in\mathbb{R}$ is the fluid pressure, with time $t\geq 0$ and space $\mathbf{z}:=(x\times y)\in\Omega\subset\mathbb{R}^{2}$. The density and dynamic viscosity of the fluid are denoted by $\rho$ and $\mu$, respectively.
@@
$\mathbf{u}$ $=\left[\frac{4Uy(0.41-y)}{0.41^{2}},0\right]^{\top},$ $\mathbf{z}\in\Gamma_{in},$
$\mu\left(\nabla\mathbf{u}+(\nabla\mathbf{u})^{\top}\right)\cdot\mathbf{n}$ $=p\mathbf{n},$ $\mathbf{z}\in\Gamma_{out},$
$\mathbf{u}$ $=0,$ $\mathbf{z}\in\Gamma_{cyl}\cup\Gamma_{top}\cup\Gamma_{bottom},$
@@
The domain $\Omega_{fluid}$ is discretized with tetrahedral mesh elements amounting to a total of $37,514$ grid nodes. The Navier-Stokes equations are numerically solved using the FEniCS package [42]. More specifically, our solver employs the finite element method and an incremental pressure correction scheme to handle the pressure-velocity coupling. For further details about the implementation, please refer to [32]. We fix the fluid density to a value of $\rho=1\ kg/m^{3}$ and can freely vary the dynamic viscosity $\mu$, resulting in a variable Reynolds number between $[80,200]$. In this work, we focus on constructing an actively-learned ROM for the fluid velocity magnitude over the complete domain $\Omega_{fluid}$.
@@
To begin the active learning procedure, $81$ evenly distributed $Re$ values are considered from $[80,200]$. Out of these, five $Re$ values, i.e., $\{80,110,140,170,200\}$, are initially taken in the parameter set $P$, whereas, the remaining $76$ $Re$ values are placed in the candidate set $P^{*}$. Furthermore, parameter-specific snapshot matrices for the fluid magnitude are constructed by querying the FOM solver corresponding to each of the $Re$ values from $P$. Similar to the shallow water equations example, parameter-specific subspaces are then formed, retaining $99.9999\%$ of the total energy, which corresponds to $\eta=10^{-6}$.
With this problem, we intend to demonstrate the applicability of Algorithm 3.1 and terminate the active learning procedure as soon as a specific user-defined computational budget for FOM queries is satisfied. Moreover, the SDE-ActLearn-POD-NN ROM is constructed as outlined in Algorithm 4.7, where a deep NN is used to learn the POD coefficients. Once the initial set of parameter-specific subspaces are formed, the distances between all the neighboring subspaces are evaluated and the parameter pair with the most distant subspaces is identified. This is followed by the selection of a new sample from the candidate set $P^{*}$ as outlined in Algorithm 3.1. In an iterative fashion, this procedure is continued, by updating the neighboring parameter pairs, evaluating the distance between new pairs, and selecting new samples from $P^{*}$ in parameter regions surrounded by subspaces that are farthest away.
@@
Table 5.2 provides runtime for the FOM solver and the SDE-ActLearn-POD-NN ROM. The FOM solver is executed on a node of the compute cluster Mechthild available at the Max Planck Institute Magdeburg. The node is equipped with two Intel® Xeon Silver 4110 central processing units and $192$ GB of RAM. The deep NN used for this example is trained on an NVIDIA Tesla P100 graphics processing unit associated with a similar cluster node. Like the shallow water equations example, we report the time required for creating the parameter-specific subspaces along with all the distance evaluations during the active learning process under SDE-AL in Table 5.2. This amounts to $59.8689$ seconds. The runtime reported under ‘Remaining’ corresponds to the global reduced basis construction in eqs. 4.11 and 4.12, NN data preparation, as well as the NN model construction and training. Out of this, the network training takes $2072.5892$ seconds. Overall, the offline phase cost is dominated by computations towards the FOM queries. The SDE-AL framework enables selection of critical parameter locations in the offline phase, thereby, rendering an effective ROM construction. In comparison with a FOM solver query, the ROM’s online query time of $0.0043$ seconds provides us a computational speedup of greater than $\mathcal{O}(10^{6}).$